OpenStudy (anonymous):

The lines 2x - 3y = 1 and 2x + 3y = 2 are intersecting.

OpenStudy (anonymous):

True or false?

OpenStudy (anonymous):

have you tried setting up a system of equations and trying to solve it. They will intersect when they have the same x and y values.

OpenStudy (anonymous):

I don't even know where to start and how to do that.

OpenStudy (anonymous):

2x-3y=1 + 2x+3y=2 ----------- 4x =3 x=3/4

OpenStudy (anonymous):

then substitute 3/4 for x in any of the 2 equations and solve for y, if you get a real value then they intersect at those x and y values, if not then they don't intersect.

OpenStudy (anonymous):

Ohh okay! Got it! Thanks!

OpenStudy (anonymous):

sure